A random sample of 64 bags of white cheddar popcorn weighed, on avarage, 5.23 ounces with a standard deviation of 0.24 ounces. Test the hypothesis that mean = 5.5 ounces against the alternative hypothesis, mean < 5.5 ounces at the 0.05 level of significance.
1
Expert's answer
2021-09-30T13:12:05-0400

"n=64 \\

\bar{x}=5.23 \\

s= 0.24 \\

H_0: \mu = 5.5 \\

H_1: \mu <5.5"

Test-statistic:

"t = \frac{\bar{x}- \mu}{s / sqrt{n}} \\

t = \frac{5.23-5.5}{0.24 / \sqrt{64}} = -9 \\

α=0.05"

Critical value "t_{n-1, \u03b1} = t_{63,0.05}= -1.6694"

Since "|t| > t_{n-1, \u03b1}" we reject H0 and conclude, that "\mu <5.5" at 0.05 level of significance.